Assume that a procedure yields a binomial distribution with n trials and the probability of success for one trial is p. Use the given values of n and p to find the mean μ and standard deviation σ. Also, use the range rule of thumb to find the minimum usual value μ−2σand the maximum usual value μ+2σ.

n=1475​,

p=3/5
Answer #1

Given that, n = 1475 and p = 3/5 = 0.6

Mean and standard deviation of the Binomial distribution are,

μ = np-1475 * 0.6-885

σ = Vinpil-p)-V1475 * 0.6 * (1-0.6) 18.8149

Therefore,

μ-20 = 885-(2 * 18.8149) = 885-37.6298 847.3702 847.37

μ + 20 885 + (2 * 18.8149) 885 + 37.6298 922.6298 922.63

The minimum usual value = 847.37

The maximum uaual value = 922.63